An influenza vaccine is produced by two different companies. It is known that a vaccine

produced by company 1 is effective with probability 0:89, while a vaccine produced by

company 2 is effective with probability 0:93. We also know that company 1 supplies 40%

of the vaccines, while company 2 supplies 60% of the vaccines ordered by the government.


(i) What is the probability that a vaccine is effective, given that it was produced by company 2?


(ii) What is the probability that a randomly chosen vaccine from the government’s order

is not effective?


(iii) What is the probability that given a vaccine is not effective that it was produced by

company 1?


1
Expert's answer
2021-12-10T10:19:19-0500

Let "C_1" and "C_2" be the events that the influenza vaccine is manufactured by company 1 and 2 respectively. Also, let "E" be the event that the influenza vaccine produced is effective. "E'" is the event that the influenza vaccine produced is not effective.

The following probabilities are given,

"p(E|C_1)=0.89,\space p(E|C_2)=0.93,\space p(C_1)=0.40,\space p(C_2)=0.60"


"i)"

The probability that a vaccine is effective, given that it was produced by company 2 is given as,

"p(E|C_2)=0.93" as stated above.


"ii)"

We need to determine the probability that a randomly selected vaccine is effective. To do so, we shall apply the law of total probability as follows,

"p(E)=p(E|C_1)*p(C_1)+p(E|C_2)*p(C_2)=0.89*0.40+0.93*0.60=0.356+0.558=0.914"

The probability that a randomly selected vaccine is not effective is given as,

"p(E')=1-p(E)=1-0.914=0.086"

Therefore, the probability that a randomly selected vaccine is not effective is 0.086.


"iii)"

We determine the conditional probability, "p(C_1|E')" defined as,

"p(C_1|E')={p(C_1\cap E')\over p(E')}"

"P(C_1\cap E')=0.4\times0.11=0.044\space and \space p(E')=0.086"

Thus, "p(C_1|E')={0.044\over 0.086}=0.5116(4dp)"

Therefore, the probability that a randomly selected vaccine is produced by company 1 given it is not effective is 0.5116.